The thrust F developed by a rocket motor is given by F = mv +A(P 1 -P 2 ), where m is the mass of fuel ejected per unit time ,V is the velocity of the ejected gas ,A is the area of the nozzle,P 1 and P 2 are the pressures. Show that the equation given is dimensionally correct.

  1.      The thrust F developed by a rocket motor is given by F = mv +A(P1-P2), where m is the mass of fuel
ejected per unit time ,V is the velocity of the  ejected gas ,A is the area of the nozzle,P1 and P2 are the pressures. Show that the equation given is dimensionally correct.

Grade:11

1 Answers

Arun
5 years ago
Dear Karan
 
F = [M L T-2]
 
mv = [M T-1] [L T-1] = [M L T-2]
 
A(P1 – P2) = AP = [L2] [M L-1 T-2] = [M L T-2]
 
Hence verified